Steamer cruises 58 kilometers downstream in the same time that is cruises 40 kilometers upstream. The engines drive the steamer in calm water at a rate which is 20 km/hr more than the rate of the current. Find the rate of the current.
How do I set this up?

Answer by mananth(16946)

let current speed = x
steamer speed = x+20
upstream speed = (x+20-x)=20 mph
downstream speed = 2x+20
upstream distance = 40
downstream distance = 58
times are same
t=d/r
40/20 = 58/(2x+20)
cross multiply
40(2x+20)=58*(20)
80x+800=1160
80x=1160-800
80x=360
x=4.5 mph
m.ananth@hotmail.ca